# MeanSparse: Post-Training Robustness Enhancement Through Mean-Centered Feature Sparsification

### v1

@@ Abstract
We present a simple yet effective method to improve the robustness of Convolutional Neural Networks (CNNs) against adversarial examples by post-processing an adversarially trained model. Our technique, MeanSparse, cascades the activation functions of a trained model with novel operators that sparsify mean-centered feature vectors. This is equivalent to reducing feature variations around the mean, and we show that such reduced variations merely affect the model’s utility, yet they strongly attenuate the adversarial perturbations and decrease the attacker’s success rate. Our experiments show that, when applied to the top models in the RobustBench leaderboard, MeanSparse achieves a new robustness record of $72.08\%$ (from $71.07\%$) and $59.64\%$ (from $59.56\%$) on CIFAR-10 and ImageNet, respectively, in terms of AutoAttack accuracy. Code is available at https://github.com/SPIN-UMass/MeanSparse
@@ 1 Introduction
MeanSparse establishes new state-of-the-art robustness. Applying it to the leading robust model on CIFAR-10 increases $\ell_{\infty}$ AutoAttack accuracy from $71.07\%$ to $72.08\%$, while clean accuracy remains nearly unchanged ($93.27\%$ to $93.24\%$) [7]. On ImageNet-1K, MeanSparse boosts the AutoAttack accuracy of the third-ranked model to $59.64\%$, surpassing the previous best of $59.56\%$, according to RobustBench specifications[10, 7].
@@
  • Integrating MeanSparse into the top two robust models from the RobustBench benchmark [7] enhanced robustness, setting a new record on CIFAR-10 ($71.07\%$ to $72.08\%$). Additionally, applying MeanSparse to the third-ranked ImageNet model set a new AutoAttack accuracy record ($59.56\%$ to $59.64\%$).
@@ 2.1 Notations
We use bold lowercase letters to show vectors while the $i$-th element of vector $\boldsymbol{x}$ is represented by $x_{i}$. Two vector norm metrics are used throughout the paper, including $\ell_{0}=\|\cdot\|_{0}$ (number of non-zero elements) and $\ell_{2}=\|\cdot\|_{2}$ (Euclidean norm). The $\ell_{0}$ norm can also effectively penalize the variables with small values but its nonsmoothness limits its application. The proximal operator finds the minimizer of function $f$ in the vicinity of input vector $\boldsymbol{v}$ and plays an essential role in our intuition to design the mean-based sparsification operator.
@@
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a proper and lower semi-continuous function. The proximal operator (mapping) $\mathop{\mathrm{prox}}_{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ of $f$ at $\boldsymbol{v}$ is defined as:
$\mathop{\mathrm{prox}}_{f}(\boldsymbol{v})=\mathop{\mathrm{argmin}}_{\boldsymbol{x}}~{}f(\boldsymbol{x})+\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{v}\|_{2}^{2}$ (1)
The definition of proximal operator can be extended to nonsmooth functions such as $\ell_{0}$ norm. Assume $f(\boldsymbol{a})\triangleq\lambda\|\boldsymbol{a}\|_{0}$. Then the proximal operator can be calculated as [11]:
$\mathop{\mathrm{prox}}_{f}(\boldsymbol{v})=\mathcal{H}_{2\lambda}(\boldsymbol{v}),~{}\mathcal{H}_{\alpha}(v)=\begin{cases}v&|v|>\sqrt{\alpha}\\ 0&|v|\leq\sqrt{\alpha}\end{cases}$ (2)
@@ 3.1 Intuition from Regularized Optimization Objective
Figure 2: Mean-centered feature used in regularized optimization problem (3)
@@
The small changes around the feature mean value at the output of different layers are potentially non-robust features. Thus we start by penalizing those changes using $\ell_{0}$ norm in an arbitrary layer $l$ for a training sample $\boldsymbol{x}$ (extension to the whole dataset and all layers is straightforward). The original training objective can be formulated as:
$\mathcal{P}_{0}:~{}\boldsymbol{\theta}_{0}^{\star}=$ $\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}~{}L(\boldsymbol{\theta})+\gamma\|\overline{\boldsymbol{a}}(\boldsymbol{x};\boldsymbol{\theta})\|_{0},~{}~{}L(\boldsymbol{\theta})\triangleq\mathcal{L}(\boldsymbol{y}(\boldsymbol{x}),\boldsymbol{y}^{\star};\boldsymbol{\theta})$ (3)
@@
where $\boldsymbol{\theta}$ is the vector containing model parameters, $\gamma$ is a hyper-parameter, $\overline{\boldsymbol{a}}(\boldsymbol{x};\boldsymbol{\theta})$ is the centralized feature vector before the activation function in the regularized layer, $\boldsymbol{y}(\cdot)$ is the model output and $\boldsymbol{y}^{\star}$ is the target label corresponding to $\boldsymbol{x}$. Throughout the paper, we assume $\mu_{a}$ to be a fixed vector that is updated using an exponential filter during training, similar to the mean and variance vectors calculated in a batch normalization layer. Figure 2 represents the way mean-centered features for regularization are generated. One option to solve the optimization problem in (3) is via the penalty method.
@@
To solve problem (3), we introduce an approximate version of this problem based on the penalty method (for simplicity we use $\overline{\boldsymbol{a}}$ instead of $\overline{\boldsymbol{a}}(\boldsymbol{x};\boldsymbol{\theta})$):
@@
$\mathcal{P}_{\lambda}:~{}\boldsymbol{\theta}^{\star}_{0,\lambda},\boldsymbol{w}^{\star}=$ $\mathop{\mathrm{argmin}}_{\boldsymbol{\theta},\boldsymbol{w}}~{}L(\boldsymbol{\theta})+\gamma\|\boldsymbol{w}\|_{0}+\frac{1}{2\lambda}\|\boldsymbol{w}-\overline{\boldsymbol{a}}\|_{2}^{2}$ (4)
@@
where $\lambda$ is the penalty parameter, $\boldsymbol{w}$ is the penalty variable, and the optimization is with respect to both variables $\boldsymbol{\theta}$ and $\boldsymbol{w}$. Based on the penalty method (PM), we can conclude:
$\lim_{\lambda\rightarrow 0}\boldsymbol{\theta}_{0,\lambda}^{\star}=\boldsymbol{\theta}^{\star}_{0}$ (5)
@@
So to approximate $\boldsymbol{\theta}_{0}^{\star}$, one can solve problem $\mathcal{P}_{\lambda}$ and decrease $\lambda$ during the training. One approach to solve problem $\mathcal{P}_{\lambda}$ is to use block coordinate descent where the $\boldsymbol{\theta}$ and $\boldsymbol{w}$ are updated sequentially [28]. So we have two steps in each training iteration $k$.
Step1$\rightarrow$ Calculatign $\boldsymbol{w}_{k}$: To calculate $\boldsymbol{w}_{k}$, we should solve:
$\boldsymbol{w}_{k}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{w}}~{}\gamma\lambda\|\boldsymbol{w}\|_{0}+\frac{1}{2}\|\boldsymbol{w}-\overline{\boldsymbol{a}}_{k-1}\|_{2}^{2}$ (6)
@@
The problem in (6) matches the definition of the proximal operator in (1) and can be solved exactly by the proximal operator for the function $f(\boldsymbol{w})=\gamma\lambda\|\boldsymbol{w}\|_{0}$. Thus we have:
$\boldsymbol{w}_{k}=\mathop{\mathrm{prox}}_{f}(\boldsymbol{z}_{k-1})=\mathcal{H}_{2\lambda\gamma}(\overline{\boldsymbol{a}}_{k-1})$ (7)
@@
Step 2 $\rightarrow$ Calculating $\boldsymbol{\theta}_{k}$: To calculate $\boldsymbol{\theta}_{k}$, we have to solve the following problem:
$\boldsymbol{\theta}_{k}=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}~{}L(\boldsymbol{\theta})+\frac{1}{2\lambda}\|\boldsymbol{w}_{k}-\overline{\boldsymbol{a}}\|_{2}^{2}$ (8)
@@
while $\boldsymbol{w}_{k}$ is calculated using hard-thresholding operator as in (7). One step of (stochastic) gradient descent can be used to update $\boldsymbol{\theta}$ based on (8).
@@ 3.2 Mean-centered Feature Sparsification
Update rule (8) shows $\overline{\boldsymbol{a}}$ approaches $\mathcal{H}_{2\mu\gamma}(\overline{\boldsymbol{a}}_{k-1})$ by the second term while the weight for this term is increasing (equivalently $\mu$ is decreasing) during the training. Based on this intuition, we propose to use the $\mathcal{H}_{2\mu\gamma}(\overline{\boldsymbol{a}}_{k-1})$ in the forward propagation of the model. If we add the element-wise mean subtraction and addition in 2 into the hard-thresholding operator, we achieve the curve in the bottom of the first column in 1 assuming $\alpha\sigma_{\mathrm{ch}}=(2\mu\lambda)^{2}$ (for simplicity of notation, we use $Th$ instead). This curve represents a sparsification around the feature mean value $\mu_{a}$. In other words, we have the following element-wise operation over the input $\boldsymbol{a}^{(\text{in})}$:
@@ 3.3 MeanSparse Technique
Throughout the neural network, the input to the activation functions can be considered as 4-dimensional, represented as $N\times C\times H\times W$. To better capture the statistics in the representation, we use per-channel mean $\mu_{\mathrm{ch}}$ and variance $\sigma_{\mathrm{ch}}$ in sparsification operator. As a result, the mean and variance vectors are $C$-dimensional vectors. The sparsification operates on each feature channel separately, using the mean and variance of the corresponding channel.
@@
The complete pipeline to apply MeanSparse technique is formed by equipping the mean-centered sparsification operator with the three aforementioned modifications and adding it before all activation functions in the model with a shared $\alpha$ parameter. Figure 1 illustrates an example of integrating the MeanSparse technique into an adversarially trained model. In the first column, the upper figure represents the histogram of one channel features in the first layer of RaWideResNet trained adversarially over the CIFAR-10 dataset [31] along with the mean $\mu_{\mathrm{ch}}$ and standard deviation $\sigma_{\mathrm{ch}}$ over the complete training set. Based on this histogram, we design the bottom sparsification operator where the variations in $\alpha\sigma_{\mathrm{ch}}$-vicinity of the mean are blocked. The second and third column represents the input and output of the sparsification operator for one test image of the CIFAR-10 dataset $\boldsymbol{x}_{t}$ in sample space and histogram, respectively. The top figures are input and the bottom ones are output. The feature histogram for $\boldsymbol{x}_{t}$ is bimodal. One of the modes matched the $\mu_{\mathrm{ch}}$ which is the feature’s peak over the training set. Based on our intuition, the variations in this region are uninformative and can be blocked as we see in the bottom figures of the second and third columns.
By using the MeanSparse presented in this paper, one can enhance the state-of-the-art robust accuracy over both CIFAR-10 [21] and ImageNet [10] datasets, as evidenced by the RobustBench [7] rankings. Through various ablation studies, we also demonstrate the effectiveness of the proposed pipeline in multiple scenarios. This pipeline sheds light on a new approach to improve robustness without compromising clean accuracy.
@@ 4 Experiments
The proposed pipeline in Figure 1 represents the way one can integrate MeanSparse into different trained models. In this section, we present several experiments to demonstrate the effectiveness of the MeanSparse across different architectures and datasets and in improving the state-of-the-art robustness. The experiments were conducted using an NVIDIA A100 GPU.
@@ 4.2 Integration to State-of-the-Art Models
With the introduction of adversarial examples for deep learning models, their reliability was severely questioned [3]. However, several models have now been developed that present a high level of accuracy while also maintaining acceptable robustness against attacks imperceptible to humans [7]. In this part, we focus on the trained models with leading performance over CIFAR-10 [21] and ImageNet [10] datasets. We use the trained models and the proposed pipeline to integrate MeanSparse before all the activation functions. Then we evaluate the performance of the resulting models.
@@
Table 1 presents the results of integrating MeanSparse into RaWideResNet [31] and WideResNet [45] architectures trained on the CIFAR-10 dataset and ConvNeXt-L [24] and RaWideResNet [31] on the ImageNet-1K dataset using adversarial training. The results are provided for different values of $\alpha$ (note that the threshold equals $\alpha\times\sigma_{\mathrm{ch}}$). RaWideResNet [31] achieves the highest AA for $\ell_{\infty}$ untargeted attacks to the CIFAR-10 dataset submitted to RobustBench [7], while WideResNet [45] ranks second. The experiment results over the CIFAR-10 test set depict that for a range of $\alpha$ below $0.25$, the clean accuracy has decreased negligibly or almost unchanged while $\text{APGD}_{\text{ce}}$is increasing. Although these results may be attributed to overfitting, the result over the CIFAR-10 training set also represents a similar trend. So an important takeaway from CIFAR-10 results in Table 1 is that the variations around the mean value of representation can be considered as a capacity provided by the model for the attackers while its utilization is limited. MeanSparse requires one pass over CIFAR-10 training set to calculate the statistics. The time for this pass is 102 seconds and 95 seconds for the RaWideResNet and WideResNet architectures, respectively. The increase in $\text{APGD}_{\text{ce}}$evaluation time after MeanSparse integration is negligible, with only a $2\%$ increase in time.
@@
Results over ImageNet-1K presented in Table 1 show that similar outcomes are achieved when MeanSparse is integrated into models based on the ConvNeXt-L [24] and RaWideResNet [31] architectures adversarially trained. ConvNeXt-L [24] achieves the third-highest AA on the ImageNet-1K dataset for $\ell_{\infty}$ untargeted attacks submitted to RobustBench [7], while RaWideResNet [31] ranks eleventh. Note that throughout our experiment over the RaWideResNet [31] architecture for both CIFAR-10 and ImageNet-1k datasets, we do not apply the sparsification operator inside the Squeeze-and-Excitation module. The time for the MeanSparse pass over the ImageNet-1K training set is 89 minutes for the ConvNeXt-L architecture and 101 minutes for the RaWideResNet architecture. The increase in $\text{APGD}_{\text{ce}}$evaluation time after MeanSparse integration is also negligible, with only a $3\%$ increase in time.
One comment regarding Table 1 is that the results may not generalize well across attacks other than $\text{APGD}_{\text{ce}}$. Therefore, in the next experiment, we evaluate AA as a complete suite of attacks for the values of $\alpha$ whose results are colored in blue in Table 1.
Table 1: Clean accuracy and accuracy under A-PGD-CE attack [8] (represented by $\text{APGD}_{\text{ce}}$) for leading robust models on the CIFAR-10 and ImageNet-1k datasets (Base row) and after their integration with the MeanSparse technique (RaWideResNet [31] and WideResNet [45] are ranked 1st and 2nd, respectively, on Robustbench for CIFAR-10 dataset and ConvNeXt-L [24] and RaWideResNet [31] are ranked 3rd and 11th, respectively, on Robustbench [7]). The best result in each column is in bold and the threshold selected through the proposed pipeline is in blue.
@@
RaWideResNet [31] WideResNet [45] ConvNeXt-L [24] RaWideResNet [31]
@@
Figure LABEL:fig:bench:sub1 depicts the results for the Clean and AA metrics of the standard and MeanSparse-integrated versions of the RaWideResNet [31] and WideResNet [45] models, with $\alpha$ values set at $0.25$ and $0.20$ for the former and latter, respectively. RaWideResNet [31] currently holds the state-of-the-art performance in the RobustBench ranking [8] for $\ell_{\infty}$ untargeted attack. After post-processing with the proposed pipeline, the Clean accuracy drops slightly (from $93.27\%$ to $93.24\%$), while the AA significantly increases (from $71.07\%$ to $72.08\%$), setting a new record for robustness in terms of AA on the CIFAR-10 dataset. Similar results can be obtained for WideResNet [45]. Also, as a result of post-processing the robust model based on the WideResNet architecture, the resulting AA accuracy ($71.40\%$) exceeds this value for the robust model based on the RaWideResNet architecture ($71.04\%$), even though WideResNet has a less complex architecture compared to RaWideResNet, which is equipped with the Squeeze-and-Excitation module.
Figure LABEL:fig:bench:sub2 illustrates the clean and AA accuracy for two different models adversarially trained on the ImageNet-1k dataset. By carefully selecting the $\alpha$ value, while the Clean accuracy remains unchanged, the AA accuracy improves significantly. Although the ConvNeXt-L model [24] currently ranks third in AA accuracy for $\ell_{\infty}$ untargeted attacks, after postprocessing with the proposed pipeline, its AA accuracy rises to $59.64\%$ from $58.48\%$. This result surpasses the current first-ranked model, which is based on the Swin-L architecture and adversarially trained [24], and has an AA accuracy of $59.56\%$.

### v2

@@ Abstract
We present a simple yet effective method to improve the robustness of both Convolutional and attention-based Neural Networks against adversarial examples by post-processing an adversarially trained model. Our technique, MeanSparse, cascades the activation functions of a trained model with novel operators that sparsify mean-centered feature vectors. This is equivalent to reducing feature variations around the mean, and we show that such reduced variations merely affect the model’s utility, yet they strongly attenuate the adversarial perturbations and decrease the attacker’s success rate. Our experiments show that, when applied to the top models in the RobustBench leaderboard, MeanSparse achieves a new robustness record of $75.28\%$ (from $73.71\%$), $44.78\%$ (from $42.67\%$) and $62.12\%$ (from $59.56\%$) on CIFAR-10, CIFAR-100 and ImageNet, respectively, in terms of AutoAttack accuracy. Code: https://github.com/SPIN-UMass/MeanSparse
@@ 1 Introduction
MeanSparse establishes a new state-of-the-art (SOTA) in robustness. When applied to the leading robust models, it improves $\ell_{\infty}$ AutoAttack accuracy from $73.71\%$ to $75.28\%$ on CIFAR-10, from $42.67\%$ to $44.78\%$ on CIFAR-100, and from $59.56\%$ to $62.12\%$ on ImageNet, while maintaining nearly unchanged clean accuracy. Additionally, when applied to the top model for $\ell_{2}$ AutoAttack accuracy on CIFAR-10, MeanSparse boosts robust accuracy from $84.97\%$ to $87.28\%$. In all four cases, MeanSparse achieves the highest ranking according to RobustBench specifications [9].
@@
  • Integrating MeanSparse into the top robust models from the RobustBench benchmark [9] improved robustness, establishing new records for $\ell_{\infty}$ AutoAttack accuracy: on CIFAR-10 (from $73.71\%$ to $75.28\%$), CIFAR-100 (from $42.67\%$ to $44.78\%$), and ImageNet (from $59.56\%$ to $62.12\%$). Additionally, new $\ell_{2}$ AutoAttack accuracy record was set on CIFAR-10 (from $84.97\%$ to $87.28\%$).
@@ 2.1 Notations
We use bold lowercase letters to show vectors while the $i$-th element of vector $\bm{x}$ is represented by $x_{i}$. Two vector norm metrics are used throughout the paper, including $\ell_{0}=\|\cdot\|_{0}$ (number of non-zero elements) and $\ell_{2}=\|\cdot\|_{2}$ (Euclidean norm). The $\ell_{0}$ norm can also effectively penalize the variables with small values but its nonsmoothness limits its application. The proximal operator finds the minimizer of function $f$ in the vicinity of input vector $\bm{v}$ and plays an essential role in our intuition to design the mean-based sparsification operator.
@@
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a proper and lower semi-continuous function. The proximal operator (mapping) $\mathop{\mathrm{prox}}_{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ of $f$ at $\bm{v}$ is defined as:
$\mathop{\mathrm{prox}}_{f}(\bm{v})=\operatorname*{arg\,min}_{\bm{x}}~{}f(\bm{x})+\frac{1}{2}\|\bm{x}-\bm{v}\|_{2}^{2}$ (1)
The definition of proximal operator can be extended to nonsmooth functions such as $\ell_{0}$ norm. Assume $f(\bm{a})\triangleq\lambda\|\bm{a}\|_{0}$. Then the proximal operator can be calculated as [13]:
$\mathop{\mathrm{prox}}_{f}(\bm{v})=\mathcal{H}_{2\lambda}(\bm{v}),~{}\mathcal{H}_{\alpha}(v)=\begin{cases}v&|v|>\sqrt{\alpha}\\ 0&|v|\leq\sqrt{\alpha}\end{cases}$ (2)
@@ 3.1 Intuition from Regularized Optimization Objective
Figure 2: Mean-centered feature used in regularized optimization problem equation 3
@@
The small changes around the feature mean value at the output of different layers are potentially non-robust features. Thus we start by penalizing those changes using $\ell_{0}$ norm in an arbitrary layer $l$ for a training sample $\bm{x}$ (extension to the whole dataset and all layers is straightforward). The original training objective can be formulated as:
$\mathcal{P}_{0}:~{}\bm{\theta}_{0}^{\star}=$ $\operatorname*{arg\,min}_{\bm{\theta}}~{}L(\bm{\theta})+\gamma\|\overline{\bm{a}}(\bm{x};\bm{\theta})\|_{0},~{}~{}L(\bm{\theta})\triangleq\mathcal{L}(\bm{y}(\bm{x}),\bm{y}^{\star};\bm{\theta})$ (3)
@@
where $\bm{\theta}$ is the vector containing model parameters, $\gamma$ is a hyper-parameter, $\overline{\bm{a}}(\bm{x};\bm{\theta})$ is the centralized feature vector before the activation function in the regularized layer, $\bm{y}(\cdot)$ is the model output and $\bm{y}^{\star}$ is the target label corresponding to $\bm{x}$. Throughout the paper, we assume $\mu_{a}$ to be a fixed vector that is updated using an exponential filter during training, similar to the mean and variance vectors calculated in a batch normalization layer. Figure 2 represents the way mean-centered features for regularization are generated. One option to solve the optimization problem in equation 3 is via the penalty method.
@@
To solve problem equation 3, we introduce an approximate version of this problem based on the penalty method (for simplicity we use $\overline{\bm{a}}$ instead of $\overline{\bm{a}}(\bm{x};\bm{\theta})$):
@@
$\mathcal{P}_{\lambda}:~{}\bm{\theta}^{\star}_{0,\lambda},\bm{w}^{\star}=$ $\operatorname*{arg\,min}_{\bm{\theta},\bm{w}}~{}L(\bm{\theta})+\gamma\|\bm{w}\|_{0}+\frac{1}{2\lambda}\|\bm{w}-\overline{\bm{a}}\|_{2}^{2}$ (4)
@@
where $\lambda$ is the penalty parameter, $\bm{w}$ is the penalty variable, and the optimization is with respect to both variables $\bm{\theta}$ and $\bm{w}$. Based on the penalty method (PM), we can conclude:
$\lim_{\lambda\rightarrow 0}\bm{\theta}_{0,\lambda}^{\star}=\bm{\theta}^{\star}_{0}$ (5)
@@
So to approximate $\bm{\theta}_{0}^{\star}$, one can solve problem $\mathcal{P}_{\lambda}$ and decrease $\lambda$ during the training. One approach to solve problem $\mathcal{P}_{\lambda}$ is to use block coordinate descent where the $\bm{\theta}$ and $\bm{w}$ are updated sequentially [31]. So we have two steps in each training iteration $k$.
Step1$\rightarrow$ Calculatign $\bm{w}_{k}$: To calculate $\bm{w}_{k}$, we should solve:
$\bm{w}_{k}$ $=\operatorname*{arg\,min}_{\bm{w}}~{}\gamma\lambda\|\bm{w}\|_{0}+\frac{1}{2}\|\bm{w}-\overline{\bm{a}}_{k-1}\|_{2}^{2}$ (6)
@@
The problem in equation 6 matches the definition of the proximal operator in equation 1 and can be solved exactly by the proximal operator for the function $f(\bm{w})=\gamma\lambda\|\bm{w}\|_{0}$. Thus we have:
$\bm{w}_{k}=\mathop{\mathrm{prox}}_{f}(\bm{z}_{k-1})=\mathcal{H}_{2\lambda\gamma}(\overline{\bm{a}}_{k-1})$ (7)
@@
Step 2 $\rightarrow$ Calculating $\bm{\theta}_{k}$: To calculate $\bm{\theta}_{k}$, we have to solve the following problem:
$\bm{\theta}_{k}=\operatorname*{arg\,min}_{\bm{\theta}}~{}L(\bm{\theta})+\frac{1}{2\lambda}\|\bm{w}_{k}-\overline{\bm{a}}\|_{2}^{2}$ (8)
@@
while $\bm{w}_{k}$ is calculated using hard-thresholding operator as in equation 7. One step of (stochastic) gradient descent can be used to update $\bm{\theta}$ based on equation 8.
@@ 3.2 Mean-centered Feature Sparsification
Update rule equation 8 shows $\overline{\bm{a}}$ approaches $\mathcal{H}_{2\lambda\gamma}(\overline{\bm{a}}_{k-1})$ by the second term while the weight for this term is increasing (equivalently $\lambda$ is decreasing) during the training. Based on this intuition, we propose to use the $\mathcal{H}_{2\lambda\gamma}(\overline{\bm{a}}_{k-1})$ in the forward propagation of the model. If we add the element-wise mean subtraction and addition in 2 into the hard-thresholding operator, we achieve the curve in the bottom of the first column in 1 assuming $\alpha\sigma_{\mathrm{ch}}=(2\lambda\gamma)^{2}$ (for simplicity of notation, we use $Th$ instead). This curve represents a sparsification around the feature mean value $\mu_{a}$. In other words, we have the following element-wise operation over the input $\bm{a}^{(\text{in})}$:
@@ 3.3 MeanSparse Technique
Throughout the neural network, the input to the activation functions can be considered as 4-dimensional, represented as $N\times C\times H\times W$ ($N$, $C$, $H$ and $W$ represents the batch size, channels, height and width, respectively). To better capture the statistics in the representation, we use per-channel mean $\mu_{\mathrm{ch}}$ and variance $\sigma_{\mathrm{ch}}$ in sparsification operator. As a result, the mean and variance vectors are $C$-dimensional vectors. The sparsification operates on each feature channel separately, using the mean and variance of the corresponding channel.
@@
The complete pipeline to apply MeanSparse technique is formed by equipping the mean-centered sparsification operator with the three aforementioned modifications and adding it before all activation functions in the model with a shared $\alpha$ parameter. Figure 1 illustrates an example of integrating the MeanSparse technique into an adversarially trained model. In the first column, the upper figure represents the histogram of one channel features in the first layer of RaWideResNet trained adversarially over the CIFAR-10 dataset [34] along with the mean $\mu_{\mathrm{ch}}$ and standard deviation $\sigma_{\mathrm{ch}}$ over the complete training set. Based on this histogram, we design the bottom sparsification operator where the variations in $\alpha\sigma_{\mathrm{ch}}$-vicinity of the mean are blocked. The second and third column represents the input and output of the sparsification operator for one test image of the CIFAR-10 dataset $\bm{x}_{t}$ in sample space and histogram, respectively. The top figures are input and the bottom ones are output. The feature histogram for $\bm{x}_{t}$ is bimodal. One of the modes matched the $\mu_{\mathrm{ch}}$ which is the feature’s peak over the training set. Based on our intuition, the variations in this region are uninformative and can be blocked as we see in the bottom figures of the second and third columns.
By using the MeanSparse presented in this paper, one can enhance the SOTA robust accuracy over CIFAR-10, CIFAR-100 [23] and ImageNet [12] datasets, as evidenced by the RobustBench [9] rankings. Through various ablation studies, we also demonstrate the effectiveness of the proposed pipeline in multiple scenarios. This pipeline sheds light on a new approach to improve robustness without compromising clean accuracy.
@@ 4 Experiments
The proposed pipeline in Figure 1 represents the way one can integrate MeanSparse into different trained models. In this section, we present several experiments to demonstrate the effectiveness of the MeanSparse across different architectures and datasets and in improving the SOTA robustness. The experiments were conducted using an NVIDIA A100 GPU.
@@ 4.2 Integration to State-of-the-Art Models
With the introduction of adversarial examples for deep learning models, their reliability was severely questioned [5]. However, several models have now been developed that present a high level of accuracy while also maintaining acceptable robustness against attacks imperceptible to humans [9]. In this part, we focus on the trained models with leading performance over CIFAR-10, CIFAR-100 [23] and ImageNet [12] datasets. We use the trained models and the proposed pipeline to integrate MeanSparse before all the activation functions. Then we evaluate the performance of the resulting models.
Figure LABEL:fig:spaf shows that integrating MeanSparse into the SOTA robust model in the RobustBench ranking [10] across various benchmark datasets achieves a new SOTA in robustness. Figure LABEL:fig:bench:sub1 depicts the results for the Clean and AA metrics of the standard and MeanSparse-integrated versions of the WideResNet-94-16 [3], RaWideResNet-70-16 [34] and WideResNet-70-16 [48] over CIFAR-10 dataset. The original WideResNet-94-16 [3] currently holds the SOTA performance in the RobustBench ranking [10] for $\ell_{\infty}$ untargeted attack. After post-processing with the proposed pipeline, the Clean accuracy drops slightly (from $93.68\%$ to $93.63\%$), while the AA significantly increases (from $73.71\%$ to $75.28\%$), setting a new record for robustness in terms of AA on the CIFAR-10 dataset. Similar results can be obtained for RaWideResNet-70-16 [34] and WideResNet-70-16 [48]. Also, as a result of post-processing the robust model based on the WideResNet-70-16 architecture, the resulting AA accuracy ($71.41\%$) exceeds this value for the robust model based on the RaWideResNet architecture ($71.07\%$), even though WideResNet has a less complex architecture compared to RaWideResNet, which is equipped with the Squeeze-and-Excitation module.
Figure LABEL:fig:bench:sub2 represents the result of integrating MeanSparse into the current SOTA robust model (WideResNet-70-16 [48]) in terms of AA accuracy with $\ell_{2}$ distance metric. Similarly, while the clean accuracy is almost unchanged (from $95.54\%$ to $95.49\%$), the AA accuracy is improved (from $84.97\%$ to $87.28\%$).
Figure LABEL:fig:bench:sub3 illustrates the clean and AA accuracy for three different models adversarially trained on the ImageNet dataset (Swin-L [27], ConvNeXt-L [26] and RaWideResNet [34]). By carefully selecting the $\alpha$ value, while the Clean accuracy remains unchanged, the AA accuracy improves significantly. The Swin-L model [27] is currently the SOTA robust model over the ImageNet dataset. When integrating this model with MeanSparse, while clean accuracy drops slightly (from $78.92\%$ to $78.86\%$), the AA accuracy is increased (from $59.56\%$ to $62.12\%$) setting a new SOTA.
Figure LABEL:fig:bench:sub4 shows the results of integrating MeanSparse with the current state-of-the-art (SOTA) robust model on the CIFAR-100 dataset (WideResNet-70-16 [48]). For this model, the clean and AutoAttack (AA) accuracies are $75.22\%$ and $42.67\%$, respectively. After integrating MeanSparse, these accuracies change to $75.17\%$ and $44.78\%$, establishing a new SOTA robust model with nearly identical clean accuracy.
@@
Table 1 presents the results of integrating MeanSparse into RaWideResNet [34] and WideResNet [48] architectures trained on the CIFAR-10 dataset and ConvNeXt-L [26] and RaWideResNet [34] on the ImageNet dataset using adversarial training. The results are provided for different values of $\alpha$ (note that the threshold equals $\alpha\times\sigma_{\mathrm{ch}}$). The experiment results over the CIFAR-10 test set depict that for a range of $\alpha$ below $0.25$, the clean accuracy has decreased negligibly or almost unchanged while $\text{APGD}_{\text{ce}}$is increasing. Although these results may be attributed to overfitting, the result over the CIFAR-10 training set also represents a similar trend. So an important takeaway from CIFAR-10 results in Table 1 is that the variations around the mean value of representation can be considered as a capacity provided by the model for the attackers while its utilization is limited. MeanSparse requires one pass over CIFAR-10 training set to calculate the statistics. The time for this pass is 102 seconds and 95 seconds for the RaWideResNet and WideResNet architectures, respectively. The increase in $\text{APGD}_{\text{ce}}$evaluation time after MeanSparse integration is negligible, with only a $2\%$ increase in time.
@@
Results over ImageNet presented in Table 1 show that similar outcomes are achieved when MeanSparse is integrated into models based on the ConvNeXt-L [26] and RaWideResNet [34] architectures adversarially trained. Note that throughout our experiment over the RaWideResNet [34] architecture for both CIFAR-10 and ImageNet-1k datasets, we do not apply the sparsification operator inside the Squeeze-and-Excitation module. The time for the MeanSparse pass over the ImageNet-1K training set is 89 minutes for the ConvNeXt-L architecture and 101 minutes for the RaWideResNet architecture. The increase in $\text{APGD}_{\text{ce}}$evaluation time after MeanSparse integration is also negligible, with only a $3\%$ increase in time.
Table 1: Clean accuracy and accuracy under A-PGD-CE attack [10] (represented by $\text{APGD}_{\text{ce}}$) for leading robust models on the CIFAR-10 and ImageNet-1k datasets (Base row) and after their integration with the MeanSparse technique (RaWideResNet [34] and WideResNet [48] are leading models in Robustbench ranking for CIFAR-10 dataset and ConvNeXt-L [26] and RaWideResNet [34] are highly ranked models for ImageNet dataset in Robustbench [9]). The best result in each column is in bold and the threshold selected through the proposed pipeline is in blue.
@@
RaWideResNet WideResNet ConvNeXt-L RaWideResNet